OpenStudy (anonymous):

Which two operations would you use to solve the equation x/4 + 8 = 22? add and divide subtract and divide add and multiply subtract and multiply

OpenStudy (anonymous):

When you see an equation like this, you want to isolate the variable. In order to do this, you can think of performing opposite operations. The variable is divided by 4 and then 8 is added. What is the opposite of adding 8? What is the opposite of dividing by 4? The answer to those two questions is the answer to your question.

OpenStudy (anonymous):

Mattfeury Wait

OpenStudy (anonymous):

Help ME!!!

OpenStudy (anonymous):

IN NEED YOUR HELP!!

OpenStudy (anonymous):

What do you need?

OpenStudy (anonymous):

I just explained how to get the answer.

OpenStudy (anonymous):

Its Only For Mttfeury Sory!1 :D

OpenStudy (anonymous):

which is add and multiply right?

OpenStudy (anonymous):

Opposite.

OpenStudy (anonymous):

You need to "undo" the addition and division.

OpenStudy (anonymous):

add and divde

OpenStudy (anonymous):

The multiply is correct. What is the opposite of adding something?

OpenStudy (anonymous):

What is the opposite of adding something?

OpenStudy (anonymous):

Subtracting

OpenStudy (anonymous):

Yes, so you would subtract 8 and then multiply by 4. This would leave the x by itself, so you have effectively solved for x.

OpenStudy (anonymous):

(x/4) - 8 + 8 =22 - 8 (x/4) = 14 (x/4) * 4 = 14*4 x = 56